Asset Publisher

Herschel Status Report - October 2009

Herschel Status Report - October 2009

Report for period 23 September to 21 October 2009Mission operations of the Herschel space observatory continued nominally during the reporting period, with the spacecraft and subsystems all performing as expected. Investigations for the recovery of the HIFI instrument are progressing well, and a plan for switching on the instrument's back-up electronics is being finalized.

Spacecraft

The spacecraft is operating nominally on prime chains throughout.

Autonomous star tracker switch-over
Investigation into an issue with the star trackers is ongoing. Autonomous switch-over between the two star trackers remains disabled in a way that still allows diagnostics to be accumulated of what is happening on-board.

Payload

As of mid-October, Herschel was 70% through the Performance Verification Phase (PVP) of its three instruments. For PACS and the SPIRE photometer, PVP will likely be completed by the end of October with the exception of activities requiring Neptune as a target - this planet only becomes visible in November. HIFI remained switched off during the reporting period. In November Herschel will be conducting a mix of activities from four post-launch phases (Commissioning, Performance Verification, Science Demonstration and Routine Science Phases), because after the temporary unavailability of HIFI and the spectrometer part of SPIRE these two instruments will require re-commissioning and additional performance verification.

The cryogenic chains are working well and maintaining the PACS and SPIRE bolometers at their operating temperatures of 300 mK.

HIFI recovery
On 2 August 2009, after more than two months of smooth operations with HIFI, an anomaly occurred resulting in the instrument switching itself off. Since then, engineers and scientists have worked to determine the cause of the anomaly. The goal was to understand what had happened before taking the risk of turning on the redundant electronics chain for HIFI and resuming observations.

There now is a coherent and consistent picture of the likely sequence of events leading to the anomalous behaviour on 2 August. A chain of events led to an overload in some components in a power converter of the Local Oscillator Control Unit (LCU) on the instrument's prime electronics chain, causing the permanent failure of one of the diodes.

The investigation team's report is being drafted and the detailed HIFI recovery plan is being prepared for formal approval. The HIFI recovery scenario and timeline is currently being elaborated and, pending the formal approval of the proposed plans, it is anticipated to resume HIFI operations (re-commissioning on its redundant signal chain and completing its PVP) in the second half of November.

Ground Segment

Mission Operations
During the reporting period, mission operations have been conducted with the support of the ESA New Norcia ground station. Observational data stored on-board Herschel was received on ground during communication passes lasting approximately three hours. The ground segment is operating nominally.

Future Milestones

  • Second half of November 2009: Formal approval of HIFI recovery plan and resuming of HIFI operations.
  • March 2010: Herschel Observatory Routine Science Phase Readiness Review


---
Legal disclaimer
This report is based on the Herschel mission manager's report dated 21 October 2009. Please see the copyright section of the legal disclaimer (linked from the home page http://sci.esa.int) for terms of use.

Last Update: 1 September 2019
17-Oct-2019 03:56 UT

ShortUrl Portlet

Shortcut URL

https://sci.esa.int/s/wRe7BvA

Images And Videos

Related Publications

Related Links

See Also

Documentation